40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 13 19:36:36 2014, Koji, Summary, LSC, PRM 2nd/4th violin filter added 
    Reply  Tue Feb 25 18:16:22 2014, Jenne, Summary, LSC, Even more violin filters PRMvio2.png
    Reply  Tue Apr 1 21:45:01 2014, Jenne, Summary, LSC, PRM violin notch not at correct freq? 
Message ID: 9674     Entry time: Tue Feb 25 18:16:22 2014     In reply to: 9634
Author: Jenne 
Type: Summary 
Category: LSC 
Subject: Even more violin filters 

A new violin mode at 1303 Hz was ringing up this afternoon.  Rana and I added a notch for this.

RXA: while the mode at 1303.6 Hz was ringing down, I used the narrowband DTT technique to measure the Q (after turning on the notch in SUS-PRM_LSC). So its another frequency in the PRM (not the BS).

The time that it takes for 2 -foldings is 652 s, which implies that Q = pi*f*tau = 1.3e6. This seems too high by a factor of ~10, so my guess is that there is still some feedback path happening. The previous bandstop filter was centered around 1285 Hz and seems also weird that the PRM would have 2 violin modes with such different frequencies. Is the mirror rotated around the optic axis such that the standoffs are not at the same height?

Attachment 1: PRMvio2.png  481 kB  Uploaded Tue Feb 25 23:00:32 2014  | Hide | Hide all
PRMvio2.png
ELOG V3.1.3-